OpenStudy (anonymous):

will give medal 4 best answer

OpenStudy (ronlover101):

Ok what is your question?

OpenStudy (jazzyluvsu123):

The reciprocal does not have a negative sign. It's 1 divided by the number. When you divide a fraction by one, you flip flop. So, its positive 12/7

OpenStudy (anonymous):

whoa tel it in numbers

OpenStudy (ronlover101):

7/12=12/7 that is the answer

OpenStudy (ronlover101):

2/3 divide by 1/8 = 2/3 multiply 8/1 = (2 multiply 8) divide by (3 multiply 1) = 16 divide by 3 =5.33 ( accurate to 2 decimal places)

OpenStudy (jazzyluvsu123):